OpenStudy (anonymous):

A college has a student to teacher ratio of 25 to 2. If there are 178 teachers at the college, how many students attend the college?

OpenStudy (anonymous):

divide the amount of teachers by 2 and then multiply that number by 25

jimthompson5910 (jim_thompson5910):

let x = # of students student/teacher = student/teacher 25/2 = x/178 25*178 = 2*x I'll let you finish up

OpenStudy (anonymous):

so i think 2225

jimthompson5910 (jim_thompson5910):

you are correct